OpenStudy (eva12):

can someone please help me i have a test tomorrow on this, which of the following equations is linear? select all that apply. A)(4x-3y)/19=(-3x+2y)/11 B)x^2+y^2=5 C)4x+7y=13 D)sqrt(x)=y E)sinx+8y=5 F)(4x-3y)/19=11/(-3x+2y) G)4x=7y+13

OpenStudy (anonymous):

a linear function can be represented on this form: f(x)=ab+b

OpenStudy (anonymous):

f(x)=ax+b

OpenStudy (anonymous):

2 og the answers are correct

OpenStudy (anonymous):

you've got your 2 variables x and y=f(x) and your constant b

OpenStudy (eva12):

A,C,E,F are linear?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

A,C,G are linear

OpenStudy (anonymous):

^the final equation should be in y=m*x+c form to be a linear equation

OpenStudy (eva12):

thank very much

OpenStudy (anonymous):

ye a.c.g